# arXiv:2602.24042v1[cs.DS] 27 Feb 2026

### v1

@@ 1 Introduction
For $0<\varepsilon\ll 1$, consider an SK instance with $n\gg\frac{1}{\varepsilon}$ identical items, each item $i$ with value $v_{i}=\varepsilon$ and size $S_{i}\sim\operatorname{Ber}(\varepsilon)$ (i.e., $S_{i}=1$ with probability $\varepsilon$, and 0 otherwise). The expected item size is $\varepsilon$. Normalize the knapsack capacity to $1$.
@@
    $\operatorname{Non-Risky-SK}$: Expected value $2-O(\varepsilon)$, overflows with probability $1$.
@@ 1.1 Our Results
(**) [FLX18] give a fully adaptive policy that is an $8$-approximation to the optimal adaptive policy. We improve upon their approximation ratio even though our policy uses only constant-many adaptive choices.
@@
In Section 5, for any small constant $\varepsilon\in(0,1)$, we show that with $k=\tilde{O}(1/\varepsilon)$ adaptive queries, the $k$-$n$ semi-adaptivity gap is upper-bounded by $2e+1+\sqrt{\varepsilon}\approx 6.44+\sqrt{\varepsilon}$. The previously best known fully adaptive policy for $\operatorname{Risky-SK}$ provides an $8$-approximation [FLX18]. Thus, we achieve an improvement both in the approximation ratio and in the number of adaptive choices. We also deduce an upper-bound for $\operatorname{Non-Risky-SK}$. These results demonstrate that a constant number of adaptive choices suffice to obtain an additional significant improvement.
@@
  • For instances consisting only of small items, while we cannot use the simple greedy non-adaptive policy of [DGV08] that obtains an approximation of $2+O(\varepsilon)$ for $\operatorname{Non-Risky-SK}$, we design a $k$-semi-adaptive policy (Algorithm 3). We show this policy achieves approximations of $8$, $6.75$ and $2e+O(\sqrt{\varepsilon})$ for $k=0$, $k=1$ and $k=\Theta\left(1/\sqrt{\varepsilon}\right)$, respectively (2), and of $2\left(\nicefrac{{k+2}}{{k+1}}\right)^{k+2}+O(k\varepsilon)$ for general $k$ (5). This demonstrates a smooth trade-off between the number of adaptive choices and resulting approximation, as illustrated in Fig. 5.
@@
The power of a single adaptive choice is of course limited: We prove a lower bound of $\approx 1.18$ on the $1\text{-}n$ semi-adaptivity gap by first strengthening the lower bound on the standard adaptivity gap from $1.5$ ([LV14]) to $2$, then applying Lemma 1 that connects the full and semi-adaptivity gaps. The strengthened lower bound is again established through analysis of a simple class of instances, those with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions. This natural class generalizes Example 1, and is a generalization of a class of Bernoulli instances studied by [LV14].
@@
We summarize the improvements to previously known approximation bounds, achieved via the techniques we develop for the semi-adaptivity gap analyses, and displayed in Tables 1 and 2: For $\operatorname{Risky-SK}$, we first improve the best known adaptivity gap upper and lower bounds. We remark that our upper bound applies to the ordered adaptive SK model of [DGV08] as well. [^2] In this model, the policy must process the items in some given order. As [DGV08] mention, any non-adaptive policy for $\operatorname{Risky-SK}$ with an adaptivity gap of $G$ has the same adaptivity gap for the ordered adaptive model. Therefore, 1 applies to this model as well. Our algorithm (Algorithm 1) is also better in terms of running time, running in $O(n\log n)$ compared to the $O(\text{poly}(n))$ time of [DGV08]. Second, our semi-adaptive policy that achieves a $\tilde{O}(1/\varepsilon)\text{-}n$ semi-adaptivity gap of $6.44$ improves upon the best known $8$-approximate fully adaptive policy of [FLX18]. The improvement is both in the approximation ratio and in the number of adaptive choices. Third, for $\operatorname{Non-Risky-SK}$, we strengthen the lower bound on the previously-known adaptivity gap, and improve the upper bound for the special case of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions.
@@ 1.2 Related Work
For $\operatorname{Risky-SK}$[DGV08] give a non-adaptive policy with an adaptivity gap of $9.5$ (which we improve). [LV14] show that the adaptivity gap is exactly $1.5$ for instances with two items, and provide an alternative upper bound of $11.66$ for general instances, as well as an upper bound of $8$ for Bernoulli size distributions. [FLX18] develop an $(8+\varepsilon)$-approximate fully adaptive policy. We improve on this result in two respects: we provide a $2e+1\approx 6.44$-approximate policy, and it is only $\tilde{O}(1/\varepsilon)$-semi-adaptive.
@@ 2 Preliminaries
A stochastic knapsack instance $I$ of size $n$ consists of $n$ items ($|I|=n$). Each item $i\in[n]$ is represented by a pair $(v_{i},F_{i})$, where $v_{i}\geq 0$ is a (deterministic) value, and $F_{i}$ is a distribution from which a random size $S_{i}$ is independently drawn.
The goal is to insert items into a knapsack of capacity $1$ (normalized w.l.o.g.), maximizing the total value in expectation (where the expectation is taken over the randomness of the item sizes). If the total size of the items inserted exceeds $1$, we say that a knapsack overflow has occurred, and no additional item can be inserted. There are two variants of the problem: In $\operatorname{Non-Risky-SK}$, a knapsack overflow results in losing only the value of the overflowing item, while in $\operatorname{Risky-SK}$, an overflow results in losing the entire accumulated value of items inserted into the knapsack, leading to a total value of 0.
@@
For any $k$-semi-adaptive policy $p$, let $\text{val}(p,I)$ be the value obtained by policy $p$ executed on instance $I$. Let $A_{k}(I):=\max_{p\in\mathscr{P}_{k}}\operatorname{\mathbb{E}}\left[\text{val}(p,I)\right]$, [^3] As mentioned by [DGV08], we use $\max$ since one can show that the supremum implicit in the definition of $A_{k}(I)$ is attained. where $\mathscr{P}_{k}$ is the set of all $k$-semi adaptive policies. Let $\operatorname{ADAPT}(I):=A_{n}(I)$ and $\operatorname{ALG}(I):=A_{0}(I)$ be the expected value of the optimal adaptive and non-adaptive policies, respectively. For these definitions, we do not restrict the computational complexity of computing these policies, although all of our positive guarantees are achieved by polytime algorithms. We also note that randomization does not yield a better approximation ratio: In any state (defined by the accumulated reward, remaining items and the remaining unoccupied knapsack size), the optimal strategy is deterministic.
@@
For a stochastic knapsack instance $I$, the adaptivity gap is defined as $G(I):=\operatorname{ADAPT}(I)/\operatorname{ALG}(I)$. For a family of instances $\mathcal{I}$, $G(\mathcal{I}):=\sup_{I\in\mathcal{I}}G(I)$. The adaptivity gap of the SK problem is the supremum over all instances $G:=\sup_{I}G(I)$.
@@
For any stochastic knapsack instance $I$,
@@ 2.1 Auxiliary Definitions
In words, the effective value of an item is an upper bound on the value any policy may obtain by inserting the item. Relatedly, the mean truncated size of an item $i$ is the mean size of the item where any size realization bigger than $1$ is relaxed to $1$.
@@
The mean truncated size of item $i$ is defined as $\mu_{i}:=\mathbb{E}[\min(S_{i},1)]$. For a set $A$ of items, define $\mu(A):=\sum_{i\in A}\mu_{i}$.
@@ 3.2 The Power of a Single Adaptive Choice: 0-11 and 11-$n$ semi-adaptivity gaps (Sections 6-7)
Computing the supremum of the gaps over all instances is a hard problem. Indeed, so far the only existing lower bounds for $\operatorname{Risky-SK}$ [LV14] and $\operatorname{Non-Risky-SK}$ [DGV08] are obtained via analyzing a simple $2$- or $3$-item instance. We introduce a $3$-step approach:
@@
We apply this approach multiple times. In our $0\text{-}1$ semi-adaptivity gap analysis (Section 6), each of the three steps preserves the adaptivity gap, allowing us to derive a tight result—that is, we obtain both the upper and lower bounds simultaneously. In Section 7, we apply the same three-step method with a recursive construction: starting from a “worst-case” instance (i.e., one with the largest gap) where the optimal adaptive decision tree has height k1k-1, we construct a new instance with height $k$. This leads to a recursive optimization that converges to a gap of $2$.
@@
We apply the same three-step framework, now extended from $\mathcal{H}_{2}$ to more general $\mathcal{H}_{k}$ instances. We begin by establishing a lower bound of $2$ on the adaptivity gap for $\operatorname{Risky-SK}$. By focusing on simpler $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions (see Definition 13), we significantly reduce the number of viable options for the optimal non-adaptive policy to linear in $n$ (Step 1).
In Step 2, we equalize (or restrict) the expected value across different non-adaptive choices. This leads to a recursive formula that bounds the adaptivity gap $G_{k}$ for an instance $I_{k}\in\mathcal{H}_{k}$ in terms of $G_{k-1}$, the gap for $\mathcal{H}_{k-1}$ instances. In the final step, we show that this recursive bound converges to $2$ as $k\to\infty$ (8). This lower bound of $2$, combined with 1 and 7, implies a $1\text{-}n$ semi-adaptivity gap lower bound of approximately $1.18$ (3).
A similar structure underlies the upper bound of $2$ on the adaptivity gap for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ instances in the $\operatorname{Non-Risky-SK}$ setting (5). The key difference lies in Step 1, where we reduce the analysis to a carefully constructed instance that includes terminal items and admits a simple optimal adaptive decision tree (14).
@@ 4 Power of Single Adaptive Choice: 0-$n$ and 11-$n$ Gap Upper Bounds
The idea in the proof of 2 is to first show that $\operatorname{\mathbb{E}}[\mu(A)]\leq 2$ where $A$ is the random set of items an adaptive policy attempts to insert (which follows from a simple martingale property and the fact that the sum of ”truncated” inserted sizes in any execution path is at most $2$) and then use this property to show that a feasible solution to $\Phi(2)$ upper bounds $\operatorname{ADAPT}$.
@@ 4.1 An Improved Non-Adaptive Policy
By 2, if we could choose to insert a prefix block of the greedy ordering with an expected size of $\tfrac{1}{2}$ we would get an adaptivity gap of $8$: By Markov inequality w.p. at least $\nicefrac{{1}}{{2}}$ the entire block would fit the knapsack. In case the entire block fits, a value of at least $\nicefrac{{1}}{{2}}\cdot\Phi(1)$ is obtained (2), which implies an expected value of at least $\nicefrac{{1}}{{4}}\cdot\Phi(1)$. Due to 2, this is at least $\nicefrac{{1}}{{8}}\cdot ADAPT$.
@@
If $\alpha=0.5$ then the option of inserting only $B$ yields at least $\alpha(1-\alpha)\Phi(1)=0.25\ \Phi(1)\geq\frac{1}{8}ADAPT$, where the last inequality holds due to 2. Otherwise, $\alpha<0.5\leq\beta$, and thus there is such item $l$ such that $\mu_{l}=\beta-\alpha$. If $\beta>1$ then $\mu_{\ell}\geq 0.5$. By 2, $w(B)+w_{\ell}\geq\Phi(1)$, so $\max\{w(B),w_{\ell}\}\geq\tfrac{1}{2}\,\Phi(1)$. Consequently, the chosen option earns at least $\tfrac{1}{2}\,\Phi(1)$, yielding an adaptivity gap of at most $4$. Hence we may assume $\beta\in[0.5,1]$ for the remainder of the proof.
@@ 4.2 Improved Approximation Via a Single Adaptive Choice
We show an efficient algorithm (Algorithm 2) that finds a $1$-semi-adaptive policy — a policy that makes at most a single adaptive choice that has improved performance compared to the best known non-adaptive policy.
@@
Algorithm 2 computes the maximum of the term $o_{5}$ in line $7$, and the maximizing $t$ in line $17$. The term is a continuous function in $t$ and thus such a maximum exists in the closed segment $\left[\zeta,\ 1\right]$ by the Weierstrass extreme value theorem. In any case, for the analysis we give of the $8.26$ approximation it is enough for the algorithm to choose the fixed value of $t=1-\dfrac{2(\alpha-p)}{1-p}$. The reason for maximizing over $t\in\left[\zeta,\ 1\right]$ is that it may lead to a better outcome, as explained in 4.
@@
Just like in the proof of 1, we may assume that $0<p\leq\alpha<0.5\leq\beta\leq 1$, and $\gamma\in[0,0.25]$ since otherwise we get an approximation ratio of at most $8$.
@@
A computer assisted grid search suggests a slightly better lower bound of $0.25$ on $T$, yielding a slight improvement in the approximation ratio to $8$ rather than $8.26$. Note however, that our analytic bound of $T\geq 0.24215$ is almost tight: by choosing $\alpha=p=0,\beta=0.5,\gamma=0$ we get $T\leq 0.25$.

### v2

@@ 1 Introduction
For $0<\varepsilon\ll 1$, consider an SK instance with $n\gg\frac{1}{\varepsilon}$ identical items, each item $i$ with value $v_{i}=\varepsilon$ and size $S_{i}\sim\operatorname{Ber}(\varepsilon)$ (i.e., $S_{i}=1$ with probability $\varepsilon$, and 0 otherwise). The expected item size is $\varepsilon$. Normalize the knapsack capacity to 11.
@@
    $\operatorname{Non-Risky-SK}$: Expected value $2-O(\varepsilon)$, overflows with probability 11.
@@ 1.1 Our Results
(**) [FLX18] give a fully adaptive policy that is an 88-approximation to the optimal adaptive policy. We improve upon their approximation ratio even though our policy uses only constant-many adaptive choices.
@@
In Section 5, for any small constant $\varepsilon\in(0,1)$, we show that with $k=\tilde{O}(1/\varepsilon)$ adaptive queries, the $k$-$n$ semi-adaptivity gap is upper-bounded by $2e+1+\sqrt{\varepsilon}\approx 6.44+\sqrt{\varepsilon}$. The previously best known fully adaptive policy for $\operatorname{Risky-SK}$ provides an 88-approximation [FLX18]. Thus, we achieve an improvement both in the approximation ratio and in the number of adaptive choices. We also deduce an upper-bound for $\operatorname{Non-Risky-SK}$. These results demonstrate that a constant number of adaptive choices suffice to obtain an additional significant improvement.
@@
  • For instances consisting only of small items, while we cannot use the simple greedy non-adaptive policy of [DGV08] that obtains an approximation of $2+O(\varepsilon)$ for $\operatorname{Non-Risky-SK}$, we design a $k$-semi-adaptive policy (Algorithm 3). We show this policy achieves approximations of 88, $6.75$ and $2e+O(\sqrt{\varepsilon})$ for $k=0$, $k=1$ and $k=\Theta\left(1/\sqrt{\varepsilon}\right)$, respectively (2), and of $2\left(\nicefrac{{k+2}}{{k+1}}\right)^{k+2}+O(k\varepsilon)$ for general $k$ (5). This demonstrates a smooth trade-off between the number of adaptive choices and resulting approximation, as illustrated in Fig. 5.
@@
The power of a single adaptive choice is of course limited: We prove a lower bound of $\approx 1.18$ on the $1\text{-}n$ semi-adaptivity gap by first strengthening the lower bound on the standard adaptivity gap from $1.5$ ([LV14]) to 22, then applying Lemma 1 that connects the full and semi-adaptivity gaps. The strengthened lower bound is again established through analysis of a simple class of instances, those with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions. This natural class generalizes Example 1, and is a generalization of a class of Bernoulli instances studied by [LV14].
@@
We summarize the improvements to previously known approximation bounds, achieved via the techniques we develop for the semi-adaptivity gap analyses, and displayed in Tables 1 and 2: For $\operatorname{Risky-SK}$, we first improve the best known adaptivity gap upper and lower bounds. We remark that our upper bound applies to the ordered adaptive SK model of [DGV08] as well. [^2] In this model, the policy must process the items in some given order. As [DGV08] mention, any non-adaptive policy for $\operatorname{Risky-SK}$ with an adaptivity gap of $G$ has the same adaptivity gap for the ordered adaptive model. Therefore, 1 applies to this model as well. Our algorithm (Algorithm 1) is also better in terms of running time, running in $O(n\log n)$ compared to the $O(\text{poly}(n))$ time of [DGV08]. Second, our semi-adaptive policy that achieves a $\tilde{O}(1/\varepsilon)\text{-}n$ semi-adaptivity gap of $6.44$ improves upon the best known 88-approximate fully adaptive policy of [FLX18]. The improvement is both in the approximation ratio and in the number of adaptive choices. Third, for $\operatorname{Non-Risky-SK}$, we strengthen the lower bound on the previously-known adaptivity gap, and improve the upper bound for the special case of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions.
@@ 1.2 Related Work
For $\operatorname{Risky-SK}$[DGV08] give a non-adaptive policy with an adaptivity gap of $9.5$ (which we improve). [LV14] show that the adaptivity gap is exactly $1.5$ for instances with two items, and provide an alternative upper bound of $11.66$ for general instances, as well as an upper bound of 88 for Bernoulli size distributions. [FLX18] develop an $(8+\varepsilon)$-approximate fully adaptive policy. We improve on this result in two respects: we provide a $2e+1\approx 6.44$-approximate policy, and it is only $\tilde{O}(1/\varepsilon)$-semi-adaptive.
@@ 2 Preliminaries
A stochastic knapsack instance II of size $n$ consists of $n$ items ($|I|=n$). Each item $i\in[n]$ is represented by a pair $(v_{i},F_{i})$, where $v_{i}\geq 0$ is a (deterministic) value, and $F_{i}$ is a distribution from which a random size $S_{i}$ is independently drawn.
The goal is to insert items into a knapsack of capacity 11 (normalized w.l.o.g.), maximizing the total value in expectation (where the expectation is taken over the randomness of the item sizes). If the total size of the items inserted exceeds 11, we say that a knapsack overflow has occurred, and no additional item can be inserted. There are two variants of the problem: In $\operatorname{Non-Risky-SK}$, a knapsack overflow results in losing only the value of the overflowing item, while in $\operatorname{Risky-SK}$, an overflow results in losing the entire accumulated value of items inserted into the knapsack, leading to a total value of 0.
@@
For any $k$-semi-adaptive policy $p$, let $\text{val}(p,I)$ be the value obtained by policy $p$ executed on instance II. Let $A_{k}(I):=\max_{p\in\mathscr{P}_{k}}\operatorname{\mathbb{E}}\left[\text{val}(p,I)\right]$, [^3] As mentioned by [DGV08], we use $\max$ since one can show that the supremum implicit in the definition of $A_{k}(I)$ is attained. where $\mathscr{P}_{k}$ is the set of all $k$-semi adaptive policies. Let $\operatorname{ADAPT}(I):=A_{n}(I)$ and $\operatorname{ALG}(I):=A_{0}(I)$ be the expected value of the optimal adaptive and non-adaptive policies, respectively. For these definitions, we do not restrict the computational complexity of computing these policies, although all of our positive guarantees are achieved by polytime algorithms. We also note that randomization does not yield a better approximation ratio: In any state (defined by the accumulated reward, remaining items and the remaining unoccupied knapsack size), the optimal strategy is deterministic.
@@
For a stochastic knapsack instance II, the adaptivity gap is defined as $G(I):=\operatorname{ADAPT}(I)/\operatorname{ALG}(I)$. For a family of instances $\mathcal{I}$, $G(\mathcal{I}):=\sup_{I\in\mathcal{I}}G(I)$. The adaptivity gap of the SK problem is the supremum over all instances $G:=\sup_{I}G(I)$.
@@
For any stochastic knapsack instance II,
@@ 2.1 Auxiliary Definitions
In words, the effective value of an item is an upper bound on the value any policy may obtain by inserting the item. Relatedly, the mean truncated size of an item $i$ is the mean size of the item where any size realization bigger than 11 is relaxed to 11.
@@
The mean truncated size of item $i$ is defined as $\mu_{i}:=\mathbb{E}[\min(S_{i},1)]$. For a set AA of items, define $\mu(A):=\sum_{i\in A}\mu_{i}$.
@@ 3.2 The Power of a Single Adaptive Choice: 0-11 and 11-$n$ semi-adaptivity gaps (Sections 6-7)
Computing the supremum of the gaps over all instances is a hard problem. Indeed, so far the only existing lower bounds for $\operatorname{Risky-SK}$ [LV14] and $\operatorname{Non-Risky-SK}$ [DGV08] are obtained via analyzing a simple 22- or 33-item instance. We introduce a 33-step approach:
@@
We apply this approach multiple times. In our $0\text{-}1$ semi-adaptivity gap analysis (Section 6), each of the three steps preserves the adaptivity gap, allowing us to derive a tight result—that is, we obtain both the upper and lower bounds simultaneously. In Section 7, we apply the same three-step method with a recursive construction: starting from a “worst-case” instance (i.e., one with the largest gap) where the optimal adaptive decision tree has height k1k-1, we construct a new instance with height $k$. This leads to a recursive optimization that converges to a gap of 22.
@@
We apply the same three-step framework, now extended from $\mathcal{H}_{2}$ to more general $\mathcal{H}_{k}$ instances. We begin by establishing a lower bound of 22 on the adaptivity gap for $\operatorname{Risky-SK}$. By focusing on simpler $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions (see Definition 13), we significantly reduce the number of viable options for the optimal non-adaptive policy to linear in $n$ (Step 1).
In Step 2, we equalize (or restrict) the expected value across different non-adaptive choices. This leads to a recursive formula that bounds the adaptivity gap $G_{k}$ for an instance $I_{k}\in\mathcal{H}_{k}$ in terms of $G_{k-1}$, the gap for $\mathcal{H}_{k-1}$ instances. In the final step, we show that this recursive bound converges to 22 as $k\to\infty$ (8). This lower bound of 22, combined with 1 and 7, implies a $1\text{-}n$ semi-adaptivity gap lower bound of approximately $1.18$ (3).
A similar structure underlies the upper bound of 22 on the adaptivity gap for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ instances in the $\operatorname{Non-Risky-SK}$ setting (5). The key difference lies in Step 1, where we reduce the analysis to a carefully constructed instance that includes terminal items and admits a simple optimal adaptive decision tree (14).
@@ 4 Power of Single Adaptive Choice: 0-$n$ and 11-$n$ Gap Upper Bounds
The idea in the proof of 2 is to first show that $\operatorname{\mathbb{E}}[\mu(A)]\leq 2$ where AA is the random set of items an adaptive policy attempts to insert (which follows from a simple martingale property and the fact that the sum of ”truncated” inserted sizes in any execution path is at most 22) and then use this property to show that a feasible solution to $\Phi(2)$ upper bounds $\operatorname{ADAPT}$.
@@ 4.1 An Improved Non-Adaptive Policy
By 2, if we could choose to insert a prefix block of the greedy ordering with an expected size of $\tfrac{1}{2}$ we would get an adaptivity gap of 88: By Markov inequality w.p. at least $\nicefrac{{1}}{{2}}$ the entire block would fit the knapsack. In case the entire block fits, a value of at least $\nicefrac{{1}}{{2}}\cdot\Phi(1)$ is obtained (2), which implies an expected value of at least $\nicefrac{{1}}{{4}}\cdot\Phi(1)$. Due to 2, this is at least $\nicefrac{{1}}{{8}}\cdot ADAPT$.
@@
If $\alpha=0.5$ then the option of inserting only $B$ yields at least $\alpha(1-\alpha)\Phi(1)=0.25\ \Phi(1)\geq\frac{1}{8}ADAPT$, where the last inequality holds due to 2. Otherwise, $\alpha<0.5\leq\beta$, and thus there is such item $l$ such that $\mu_{l}=\beta-\alpha$. If $\beta>1$ then $\mu_{\ell}\geq 0.5$. By 2, $w(B)+w_{\ell}\geq\Phi(1)$, so $\max\{w(B),w_{\ell}\}\geq\tfrac{1}{2}\,\Phi(1)$. Consequently, the chosen option earns at least $\tfrac{1}{2}\,\Phi(1)$, yielding an adaptivity gap of at most 44. Hence we may assume $\beta\in[0.5,1]$ for the remainder of the proof.
@@ 4.2 Improved Approximation Via a Single Adaptive Choice
We show an efficient algorithm (Algorithm 2) that finds a 11-semi-adaptive policy — a policy that makes at most a single adaptive choice that has improved performance compared to the best known non-adaptive policy.
@@
Algorithm 2 computes the maximum of the term $o_{5}$ in line 77, and the maximizing $t$ in line $17$. The term is a continuous function in $t$ and thus such a maximum exists in the closed segment $\left[\zeta,\ 1\right]$ by the Weierstrass extreme value theorem. In any case, for the analysis we give of the $8.26$ approximation it is enough for the algorithm to choose the fixed value of $t=1-\dfrac{2(\alpha-p)}{1-p}$. The reason for maximizing over $t\in\left[\zeta,\ 1\right]$ is that it may lead to a better outcome, as explained in 4.
@@
Just like in the proof of 1, we may assume that $0<p\leq\alpha<0.5\leq\beta\leq 1$, and $\gamma\in[0,0.25]$ since otherwise we get an approximation ratio of at most 88.
@@
A computer assisted grid search suggests a slightly better lower bound of $0.25$ on $T$, yielding a slight improvement in the approximation ratio to 88 rather than $8.26$. Note however, that our analytic bound of $T\geq 0.24215$ is almost tight: by choosing $\alpha=p=0,\beta=0.5,\gamma=0$ we get $T\leq 0.25$.